Let T:R^3 be defined by
T(x1,x2,x3)=( x1+x3,x3,x2-x3).
Is T invertible? If yes,find a rule for T^-1 like the one which defines T. If T is not invertible, check whether T satisfies Cayley-Hamiltion theorem.
1
Expert's answer
2019-03-20T11:47:58-0400

Solution:


Let "e_1 = (1, 0, 0)", "e_2 = (0, 1, 0)", "e_3 = (0, 0, 1)" be the natural basis of "\R^3". We have "T \left( e_1 \right) = e_1", "T \left( e_2 \right) = e_3", and "T \left( e_3 \right) = e_1 + e_2 - e_3". The matrix of this transformation with respect to the introduced basis is then

"\begin{pmatrix} 1 & 0 & 1 \\ 0 & 0 & 1 \\ 0 & 1 & -1 \end{pmatrix} \, ."

This matrix is non-degenerate (its determinant is equal to -1), therefore, T is invertible. To find the inverse, we observe that "T^{-1} \left( e_1 \right) = e_1", "T^{-1} \left( e_3 \right) = e_2", and "T^{-1} \left( e_1 + e_2 - e_3 \right) = e_3". From the last equation, we have "T^{-1} \left( e_2 \right) = e_3 - T^{-1} \left( e_1 \right) + T^{-1} \left( e_3 \right) = - e_1 + e_2 + e_3". Then, since "\left( x_1 , x_2 , x_3 \right) = x_1 e_1 + x_2 e_2 + x_3 e_3", we obtain

"T^{-1} \left( x_1 , x_2 , x_3 \right) = T^{-1} \left( x_1 e_1 + x_2 e_2 + x_3 e_3 \right) \\ {} = x_1 T^{-1} \left( e_1 \right) + x_2 T^{-1} \left( e_2 \right) + x_3 T^{-1} \left( e_3 \right) \\ {} = x_1 e_1 + x_2 \left( - e_1 + e_2 + e_3 \right) + x_3 e_2 \\ {} = \left( x_1 - x_2 , x_2 + x_3 , x_2 \right) \, ."

Answer: T is invertible, with the rule "T^{-1} \left( x_1 , x_2 , x_3 \right) = \left( x_1 - x_2 , x_2 + x_3 , x_2 \right)".